# Algorithms for Generating Small Random Samples

### v1

@@ 3 Experimental Methodology
We implemented the new algorithms from Section 2, as well as other random sampling algorithms in the open source library $\rho\mu$. See Table 1 for URLs to the libary sourcecode, project website, and in the Maven Central repository. The experiments of this paper specifically use $\rho\mu$ 4.0.0.
Table 1: Important URLs for the $\rho\mu$ open source library
@@
The code to reproduce the experiments is available on GitHub in a subdirectory of the repository for the $\rho\mu$ library itself, as is the data from my runs of the experiments. Table 3 provides the URLs to the experiment sourcecode and data.
Table 3: Reproducible results: URLs to experiment code and data
Code https://github.com/cicirello/rho-mu/examples
Data https://github.com/cicirello/rho-mu/examples/data/sample-2-and-3
@@ 4.1 Sampling Pairs of Distinct Integers
This first set of results in Table 4 explores the difference in runtime performance of the various approaches to sampling random pairs of distinct integers from the half open interval $[0,n)$. It is clear from the results that the RandomPair (Algorithm 1) of this paper dominates all of the others. Its average runtime is a low constant, and does not vary as $n$ increases, at least not to any meaningful degree. If it appears that average time is possibly decreasing as $n$ increases for this algorithm, it may be, although it is clearly negligible. As $n$ increases, the probability of the assignment statement on line 4 of Algorithm 1 decreases. The insertion sampling [2] average runtime is likewise approximately constant, though a higher constant (note that the $n=1024$ case lead to unexpectedly longer average runtime for insertion sampling, and we offer no explanation). Note that insertion sampling’s theoretical runtime is $O(k^{2})$, which is essentially constant for any fixed $k$. Pool sampling [5] and reservoir sampling [11] both have an average runtime that increases linearly in $n$. Reservoir sampling’s average runtime is much higher than pool sampling, however, because reservoir sampling also generates $(n-k)$ random integers to sample $k$ integers from a set of $n$ integers, while pool sampling only needs $k$ random integers to accomplish this. Thus, reservoir is an especially poor choice for low $k$ relative to $n$.
Table 4: Sampling pairs of distinct integers: average time (nanoseconds) with 99.9% confidence intervals
@@
$n=16$ $16.110\pm 0.021$ $24.651\pm 0.938$ $55.092\pm 0.294$ $177.547\pm 0.127$
$n=64$ $15.632\pm 0.339$ $23.838\pm 0.744$ $89.842\pm 0.196$ $633.248\pm 5.437$
$n=256$ $15.543\pm 0.124$ $23.678\pm 0.208$ $300.194\pm 1.691$ $2081.356\pm 1.574$
$n=1024$ $15.460\pm 0.557$ $31.238\pm 0.133$ $1209.842\pm 10.175$ $7931.752\pm 62.058$
In the results above, we used the version of the EnhancedRandomGenerator.nextIntPair method that allocates and returns an array for the pair of integers, which is also what the various sampling algorithm implementations do. The $\rho\mu$ library also provides a version of the EnhancedRandomGenerator.nextIntPair that returns a Java record. A Java record is immutable and potentially enables the JVM to optimize in ways that it otherwise cannot. To explore this, we ran an additional experiment comparing three cases: (a) allocating and returning a new array for the integer pair, (b) using a preallocated array passed as a parameter, and (c) allocating and returning a record instead of an array. In this experiment, rather than using JMH’s blackhole to consume the generated pairs directly (as we did in the first set of results), we simulate an operation on the pair of integers to give the JVM the opportunity to exploit the immutability property of Java records. Specifically, we do a simple sum of the integers in the pairs, and return that to JMH’s blackhole.
The results of this second experiment with sampling pairs of integers is found in Table 5. The fastest of the three options is using a Java record type for the pair of integers. It was even faster than using a preallocated array. Although the difference in time is just a little over a nanosecond per sample on average, so use whichever leads to cleaner code for your use-case.
Table 5: Sampling pairs of distinct integers: average time (nanoseconds) with 99.9% confidence intervals
@@ 4.2 Sampling Triples of Distinct Integers
Table 6: Sampling triples of distinct integers: average time (nanoseconds) with 99.9% confidence intervals
@@
Table 7: Sampling triples of distinct integers: average time (nanoseconds) with 99.9% confidence intervals

### v2

@@ 3 Experimental Methodology
We implemented the new algorithms from Section 2, as well as other random sampling algorithms in the open source library $\rho\mu$. The experiments of this paper specifically use $\rho\mu$ 4.0.0. The code to reproduce the experiments is also available on GitHub, as is the raw data from my runs of the experiments. See Table 1 for URLs to the libary sourcecode, project website, and in the Maven Central repository, as well as for the source code and raw data from the experiments.
Table 1: Important URLs for the $\rho\mu$ open source library and experiments
$\rho\mu$ library
@@
Experiments
Source/data https://github.com/cicirello/small-sample-experiments
@@ 4.1 Sampling Pairs of Distinct Integers
This first set of results in Table 3 explores the difference in runtime performance of the various approaches to sampling random pairs of distinct integers from the half open interval $[0,n)$. It is clear from the results that the RandomPair (Algorithm 1) of this paper dominates all of the others. Its average runtime is a low constant, and does not vary as $n$ increases, at least not to any meaningful degree. If it appears that average time is possibly decreasing as $n$ increases for this algorithm, it may be, although it is clearly negligible. This is because as $n$ increases, the probability of the assignment statement on line 4 of Algorithm 1 decreases. The insertion sampling [2] average runtime is likewise approximately constant, though a higher constant. Note that insertion sampling’s theoretical runtime is $O(k^{2})$, which is essentially constant for any fixed $k$. Pool sampling [5] and reservoir sampling [11] both have an average runtime that increases linearly in $n$. Reservoir sampling’s average runtime is much higher than pool sampling, however, because reservoir sampling also generates $(n-k)$ random integers to sample $k$ integers from a set of $n$ integers, while pool sampling only needs $k$ random integers to accomplish this. Thus, reservoir is an especially poor choice for low $k$ relative to $n$.
Table 3: Sampling pairs of distinct integers: average time (ns) with 99.9% confidence intervals
@@
$n=16$ $15.899\pm 0.017$ $24.125\pm 0.271$ $55.035\pm 0.096$ $178.878\pm 4.264$
$n=64$ $15.580\pm 0.128$ $23.863\pm 0.038$ $89.234\pm 2.463$ $633.099\pm 2.630$
$n=256$ $15.485\pm 0.044$ $24.175\pm 0.106$ $295.983\pm 3.101$ $2087.132\pm 25.733$
$n=1024$ $15.451\pm 0.023$ $24.153\pm 0.088$ $1199.580\pm 8.026$ $7931.636\pm 25.100$
The results discussed above used the version of the EnhancedRandomGenerator.nextIntPair method that allocates and returns an array for the pair of integers, which is also what the various sampling algorithm implementations do. The $\rho\mu$ library also provides a version of the EnhancedRandomGenerator.nextIntPair that returns a Java record. A Java record is immutable and potentially enables the JVM to optimize in ways that it otherwise cannot. To explore this, we ran an additional experiment comparing three cases: (a) allocating and returning a new array for the integer pair, (b) using a preallocated array passed as a parameter, and (c) allocating and returning a record instead of an array. In this experiment, rather than using JMH’s blackhole to consume the generated pairs directly (as we did in the first set of results), we simulate an operation on the pair of integers to give the JVM the opportunity to exploit the immutability property of Java records. Specifically, we do a simple sum of the integers in the pairs, and return that to JMH’s blackhole.
Table 4 shows the results of this second experiment. The fastest of the three options uses a Java record type for the pair of integers. It was even faster than using a preallocated array. The difference in time is just a little over a nanosecond per sample on average, so use whichever leads to cleaner code for your use-case.
Table 4: Sampling pairs of distinct integers: average time (ns) with 99.9% confidence intervals
@@ 4.2 Sampling Triples of Distinct Integers
Table 5: Sampling triples of distinct integers: average time (ns) with 99.9% confidence intervals
@@
Table 6: Sampling triples of distinct integers: average time (ns) with 99.9% confidence intervals